# Betelgeuse, Betelgeuse, Betelgeuse, Betel-buddy? Constraints on the dynamical companion to  Orionis from HST

### v1

@@ Abstract
Recently, two independent analyses have asserted that the cause of the Long Secondary Period (LSP) observed in the variability spectrum of our nearest red supergiant, Betelgeuse ( Ori), is an as-yet undetected, low-mass binary companion dubbed . In this paper, we present the results of a far-UV observational campaign using the STIS echelle spectrograph on the Hubble Space Telescope aimed at detecting spectral signatures of the companion. The four-quadrant tiling pattern and timing of the observations were optimized to isolate the companion, with observations taking place during a period of maximum angular and velocity separation between Betelgeuse and the putative companion. Spectral differencing between quadrants recovers no spectral features at the companionâs velocity in excess of the background or Betelgeuseâs chromosphere, i.e. a non-detection. Having determined that is most likely a Young Stellar Object (YSO) thanks to constraints from a complementary X-ray campaign with the Chandra X-ray Observatory in a companion paper, comparison of our data against canonical spectra from YSOs in the ULLYSES database allows us to confidently exclude masses above Mâ and companion continuum or line emission in excess of âergâs-1âcm-2âĂ -1 in the FUV (1200-1700Ă ). Future observational campaigns aware of the LSP phase are needed to place deeper constraints on the spectroscopic nature of .
@@ 1 Introduction
Alpha Orionis ( Ori), also known as Betelgeuse, is the 10th brightest star in the sky (Hoffleit & Jaschek, 1982; Hoffleit & Warren, 1987) and arguably our nearest red supergiant (RSG), with modern distance estimates ranging from 168âpc (Joyce et al., 2020) to 222âpc (Harper et al., 2017) from Earth. Interest in this star has increased significantly over the past five years due in large part to a severe brightness drop of âmag in 2019 that has come to be known as âThe Great Dimmingâ (e.g. Dupree et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Harper et al., 2020a, b; Kravchenko et al., 2021; Taniguchi et al., 2022; Cannon et al., 2023; JadlovskĂœ et al., 2024; Drevon et al., 2024). Given that Betelgeuse is an evolved massive star (, e.g. Joyce et al. 2020), this event prompted speculation that the star was close to undergoing a supernova (SN) explosion, but investigation suggests that Betelgeuseâs behavior during the Great Dimming comes from the creation of a dust cloud and a subsequent mode transition (Levesque & Massey, 2020; MontargĂšs et al., 2021; Dupree et al., 2022; MacLeod et al., 2023). These phenomena occur in the stellar and circumstellar envelope, and therefore need not be connected to any late-stage core activity that would suggest an imminent death. Current consensus favors an SN being several hundred thousand years away (e.g. Joyce et al., 2020; NeuhĂ€user et al., 2022; MacLeod et al., 2023), but tension in the literature nonetheless remains (Saio et al., 2023).
@@
At the heart of the argument for or against an imminent SN is the classification of Betelgeuseâs periodicities and their driving mechanisms (Joyce et al., 2020; MacLeod et al., 2023). Betelgeuse exhibits a number of strong periodic signals, the most prominent of which are periodicities at roughly 2100 days, 400 days, and 185 days. If the 2100-day periodicity is an intrinsic variation corresponding to Betelgeuseâs fundamental (acoustic) pressure mode (FM), this would imply a radius in excess of and place the star in its core carbon fusion stage of evolution: a short-lived phase in the starâs last few hundred to few thousand years of life (Saio et al., 2023). The consensus regarding mode identification, however, is that the 400âday periodicity is the FM (Kiss et al., 2006; Joyce et al., 2020; Dupree et al., 2022; NeuhĂ€user et al., 2022; Granzer et al., 2022; MacLeod et al., 2023; MolnĂĄr et al., 2023) which represents the slowest (lowest-frequency) form of intrinsic variability caused by acoustic pressure variation.
@@
This identification, coupled with other high-fidelity constraints, requires that Betelgeuse is undergoing He fusion in the core and therefore has âyr before its eventual SN. However, it also leaves open the question of what drives the longer (even lower-frequency), 2100âday periodicity. Long periodicities occurring at sub-FM frequencies are observed in about one third of cool, luminous stars with large, radially extended convective envelopes (e.g. Payne-Gaposchkin, 1954; Wood et al., 1999, 2004; Nicholls et al., 2009; SoszyĆski, 2007; SoszyĆski et al., 2009, 2011, 2013, 2021; Pawlak, 2021, 2023), including Red Supergiants such as Betelgeuse (e.g. Kiss et al., 2006; Yang & Jiang, 2012; Percy & Khatu, 2014; Chatys et al., 2019).
@@
These sub-FM periodicities have been given the name âLong Secondary Periods,â or LSPs. There have been many mechanisms proposed to explain LSPs in general, ranging from giant cell convection (e.g. Stothers, 1972, 2010), mode interactions (e.g. Saio et al., 2015), non-radial pulsations (e.g. Wood, 2000), and dust modulation beyond the stellar surface (e.g. Wood et al., 2004; Wood & Nicholls, 2009; Percy & Shenoy, 2023). The most salient hypothesis, however, is binarity: namely, that the long secondary period is induced by an orbital companion (e.g. Wood et al. 2004; Derekas et al. 2006; SoszyĆski 2007; SoszyĆski et al. 2021).
@@
Recently, Goldberg et al. (2024) and MacLeod et al. (2025) independently proposed the existence of a low-mass companion in a âd orbit as the explanation for the LSP observed in Betelgeuse, which is present in both the starâs lightcurve and its spectroscopic radial velocity (RV). Through a joint analysis of Betelgeuseâs brightness and RV variations and the phase offset between them compared to a sample of other stars with measured LSPs and RVs, Goldberg et al. (2024) argued that the binarity hypothesis is the best explanation for Betelgeuseâs LSP and RV signal, systematically ruling out all other proposed causes of the LSP in Betelgeuse. MacLeod et al. (2025) infer the existence of an using a long baseline of RV data as well as astrometric analysis, finding a periodic signal consistent with orbital motion on the LSP timescale in both data sets. In both of these works, the LSP is attributed to a low-mass () binary companion on a âyr, âAU, low-eccentricity orbit, at relatively low inclination relative to the observer. A companion at such a low mass and wide orbital separation would not directly affect Betelgeuseâs evolution through any substantial mass transfer, but it may interact indirectly via tides or by accreting from Betelgeuseâs stellar wind.
@@
The idea that Betelgeuse could presently be a binary system has been explored in the literature. For example, from speckle-imaging measurements Karovska et al. (1986) proposed that Betelgeuse has two high-eccentricity companions, including one at a separation of 0.06â (near that of ) but a period of 2.1 years.111This particular configuration was not confirmed in follow-up work with higher-resolution instruments (e.g. Wilson et al., 1992; Kervella et al., 2009; MontargĂšs et al., 2016), nor recovered in the lightcurve or RV signal (Joyce et al., 2020; Goldberg et al., 2024; MacLeod et al., 2023, 2025). Moreover, a companion orbiting Betelgeuse is consistent with expectations for stars in Betelgeuseâs mass class (). Betelgeuse evolved from a late-O/early-B, main-sequence (MS) star. The binary fraction of late-O/early-B stars is 95% (e.g. Sana et al., 2012; de Mink et al., 2013). Though close binaries ( AU, which would have interacted by the RSG phase) show a uniform distribution of mass ratio with an excess at , at wider separations the mass ratio companions are skewed toward small mass ratios, favoring at 10AU (Rizzuto et al., 2013; Moe & Di Stefano, 2017; Offner et al., 2023). Other systems have been confirmed to have companions with low and intermediate separations, such as a large population of Cepheids in the range with binary companions on 1-20 yr periods, identified via both RV measurements (e.g. Evans et al., 2015) and X-ray emission and UV excess (Evans et al., 2022). It is therefore unsurprising that such a massive star has a companion at âAU.
@@
Moving beyond dynamical arguments, observational follow-up informed by the companionâs orbital configuration will further constrain its nature and impact on the star and its environment. Importantly, Goldberg et al. (2024) show that the phase offset between the RV and light curves requires the companion to be behind Betelgeuse during Betelgeuseâs LSP brightness minima. The Great Dimming, which prompted many immediate follow-up observations, occurred at the brightness minimum of both the stellar pulsation and the LSP. However, at the time the companion was behind the disc of Betelgeuse, so observations then (e.g. Dupree et al., 2020, 2022; Kashyap et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Kravchenko et al., 2021; Alexeeva et al., 2021; Matthews & Dupree, 2022; Taniguchi et al., 2022) would not show any emission from the companion. Epochs corresponding to other orbital configurations, on the other hand, are ripe for further exploration.
@@
MacLeod et al. (2025) further argue that the companion could plausibly induce rotational velocities of km/s via tidal interactions, matching the measured 36 year period in the lightcurve and RV (Joyce et al., 2020; MacLeod et al., 2025) as well as observations of a dipolar ârotationalâ velocity field in spatially-resolved observations (redshifted in the SE and blueshifted in the NW, see, e.g. Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018). Fast rotation in Betelgeuse has previously been argued to indicate a recent merger, as rapid rotation of a RSG at the km/s level requires an external source of angular momentum input into the starâs outer envelope (e.g. Wheeler et al., 2017; Wheeler & Chatzopoulos, 2023; Chatzopoulos et al., 2020; Nance et al., 2018; Shiber et al., 2024).222See, however, discussion by Ma et al. (2024) on under-resolved large-scale convection as a plausible origin for the Kervella et al. (2018) velocity field, a counter-hypothesis which predicts a dipole orientation which changes stochastically and decoheres on a year-long convective overturn timescale. If the companion is instead responsible, then when combined with an orbital phase inferred from Betelgeuseâs RV, the direction of rotation (PA of 48 W. of S.; Kervella et al. 2018) would constrain the companionâs projected orbital plane and thus on-sky location relative to Betelgeuse.
@@
We obtained Directorâs Discretionary Time (DDT) during Hubble Space Telescope (HST) Cycle 32 to observe Betelgeuse and its putative companion with the Space Telescope Imaging Spectrograph (STIS) instrument in the Far-UV (FUV) in a configuration never before observed. Though MacLeod et al. (2025) argue that the chromospheric emission from Betelgeuse would prevent detection of a companion, the mass rangeâwhich extends potentially up to permitted by the high-quality STELLA RVs (Goldberg et al., 2024) and consistent with the astrometric analysis of MacLeod et al. (2025)âincludes a number of scenarios that could be detectable in both the far-UV and X-ray bands (see our companion publication in OâGrady et al. 2025).
@@ 2 Expectations for UV emission
Our observing strategy is informed by the various mass estimates and constraints for , and the most likely evolutionary possibilities for its companion. In this section we discuss the range of expectations for the companionâs spectral properties, and emphasize the importance of observing at âs peak expected offset in redshift from Betelgeuse.
@@ 2.1 Possible Companion Scenarios
A mass range of â0.5â2 Mâ enables a range of possible identities for the companion. One astrophysical object satisfying the mass condition is a low-mass, âstandardâ star. Under co-natal and coeval assumptions for the system (Dolan et al., 2016; Joyce et al., 2020; NeuhĂ€user et al., 2022), such a star must in fact be a young stellar object, or YSO, on the basis of Betelgeuseâs age of roughly 10 Myr (Tognelli et al., 2011; Baraffe et al., 2015; HaemmerlĂ© et al., 2019; Amard et al., 2019). It is possible that the companion is not co-natal, but rather has been captured due to dynamical interactions, with the system then ejected from its birth cluster. However, Betelgeuseâs proper motion has not yet been traced back to a specific cluster (though see some discussion in Harper et al., 2008; Bouy et al., 2014; Bouy & Alves, 2015). Moreover, ejections generally happen within 1 Myr of cluster formation (Oh & Kroupa, 2016), but the age of Betelgeuse and all nearby clusters and OB associations are much older than 1 Myr (e.g. Blaauw, 1964; Briceño et al., 2007, 2019), so even a dynamical ejection scenario would be close to coeval.
Other possibilities include a more exotic, compact object, such as a neutron star (NS) or white dwarf (WD). In the case of a WD, it is not possible to construct a non-interacting, two-body, coeval evolutionary scenario that combines a 10-Myr-old red supergiant with even the highest-mass WD, which has a theoretical lower age limit of 40 Myr and a practical (observed) lower age limit of 70 Myr. While more complex three-body (or more) scenarios involving the deposition of a WD into Betelgeuseâs orbit from some other system are theoretically possible, the kinematic signatures of such an event are unclear. Evolutionary timescales therefore preclude the most likely versions of a WD scenario.
Additionally, in the case of a compact object, we would expect to see a bright X-ray signature of accretion: given the orbital separation, the companion is embedded within the dusty shell surrounding Betelgeuse (e.g., Haubois et al., 2019, 2023), and would interact with Betelgeuseâs wind (Aizu, 1973; Livio et al., 1986; Blondin et al., 1990; Mukai, 2017; Maldonado et al., 2025). Typical X-ray luminosities for accreting WDs are erg s-1 (Luna et al., 2013), and for accreting NSs are erg s-1 (Yungelson et al., 2019; De et al., 2022a, b). Complementary observations with Chandra were performed within a few weeks of the HST observations and revealed no such signature, as discussed in detail in OâGrady et al. (2025).

### v2

@@ Abstract
Recently, two independent analyses have asserted that the cause of the Long Secondary Period (LSP) observed in the variability spectrum of our nearest red supergiant, Betelgeuse ( Ori), is an as-yet undetected, low-mass binary companion dubbed . In this paper, we present the results of a far-UV observational campaign using the STIS echelle spectrograph on the Hubble Space Telescope aimed at detecting spectral signatures of the companion. The four-quadrant tiling pattern and timing of the observations were optimized to isolate the companion, with observations taking place during a period of maximum angular and velocity separation between Betelgeuse and the putative companion. Spectral differencing between quadrants recovers no spectral features at the companion’s velocity in excess of the background or Betelgeuse’s chromosphere, i.e. a non-detection. Having determined that is most likely a Young Stellar Object (YSO) thanks to constraints from a complementary X-ray campaign with the Chandra X-ray Observatory in a companion paper, comparison of our data against canonical spectra from YSOs in the ULLYSES database allows us to confidently exclude masses above and companion continuum or line emission in excess of erg s-1 cm-2 Å-1 in the FUV (1200-1700Å). Future observational campaigns aware of the LSP phase are needed to place deeper constraints on the spectroscopic nature of .
@@ 1 Introduction
Alpha Orionis ( Ori), also known as Betelgeuse, is the 10th brightest star in the sky (Hoffleit & Jaschek, 1982; Hoffleit & Warren, 1987) and arguably our nearest red supergiant (RSG), with modern distance estimates ranging from 168 pc (Joyce et al., 2020) to 222 pc (Harper et al., 2017) from Earth. Interest in this star has increased significantly over the past five years due in large part to a severe brightness drop of mag in 2019 that has come to be known as “The Great Dimming” (e.g. Dupree et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; Montargès et al., 2021; Harper et al., 2020a, b; Kravchenko et al., 2021; Taniguchi et al., 2022; Cannon et al., 2023; Jadlovský et al., 2024; Drevon et al., 2024). Given that Betelgeuse is an evolved massive star (, e.g. Joyce et al. 2020), this event prompted speculation that the star was close to undergoing a supernova (SN) explosion, but investigation suggests that Betelgeuse’s behavior during the Great Dimming comes from the creation of a dust cloud and a subsequent mode transition (Levesque & Massey, 2020; Montargès et al., 2021; Dupree et al., 2022; MacLeod et al., 2023). These phenomena occur in the stellar and circumstellar envelope, and therefore need not be connected to any late-stage core activity that would suggest an imminent death. Current consensus favors an SN being several hundred thousand years away (e.g. Joyce et al., 2020; Neuhäuser et al., 2022; MacLeod et al., 2023), but tension in the literature nonetheless remains (Saio et al., 2023).
@@
At the heart of the argument for or against an imminent SN is the classification of Betelgeuse’s periodicities and their driving mechanisms (Joyce et al., 2020; MacLeod et al., 2023). Betelgeuse exhibits a number of strong periodic signals, the most prominent of which are periodicities at roughly 2100 days, 400 days, and 185 days. If the 2100-day periodicity is an intrinsic variation corresponding to Betelgeuse’s fundamental (acoustic) pressure mode (FM), this would imply a radius in excess of and place the star in its core carbon fusion stage of evolution: a short-lived phase in the star’s last few hundred to few thousand years of life (Saio et al., 2023). The consensus regarding mode identification, however, is that the 400–day periodicity is the FM (Kiss et al., 2006; Joyce et al., 2020; Dupree et al., 2022; Neuhäuser et al., 2022; Granzer et al., 2022; MacLeod et al., 2023; Molnár et al., 2023) which represents the slowest (lowest-frequency) form of intrinsic variability caused by acoustic pressure variation.
@@
This identification, coupled with other high-fidelity constraints, requires that Betelgeuse is undergoing He fusion in the core and therefore has yr before its eventual SN. However, it also leaves open the question of what drives the longer (even lower-frequency), 2100–day periodicity. Long periodicities occurring at sub-FM frequencies are observed in about one third of cool, luminous stars with large, radially extended convective envelopes (e.g. Payne-Gaposchkin, 1954; Wood et al., 1999, 2004; Nicholls et al., 2009; Soszyński, 2007; Soszyński et al., 2009, 2011, 2013, 2021; Pawlak, 2021, 2023), including Red Supergiants such as Betelgeuse (e.g. Kiss et al., 2006; Yang & Jiang, 2012; Percy & Khatu, 2014; Chatys et al., 2019).
@@
These sub-FM periodicities have been given the name “Long Secondary Periods,” or LSPs. There have been many mechanisms proposed to explain LSPs in general, ranging from giant cell convection (e.g. Stothers, 1972, 2010), mode interactions (e.g. Saio et al., 2015), non-radial pulsations (e.g. Wood, 2000), and dust modulation beyond the stellar surface (e.g. Wood et al., 2004; Wood & Nicholls, 2009; Percy & Shenoy, 2023). The most salient hypothesis, however, is binarity: namely, that the long secondary period is induced by an orbital companion (e.g. Wood et al. 2004; Derekas et al. 2006; Soszyński 2007; Soszyński et al. 2021).
@@
Recently, Goldberg et al. (2024) and MacLeod et al. (2025) independently proposed the existence of a low-mass companion in a d orbit as the explanation for the LSP observed in Betelgeuse, which is present in both the star’s lightcurve and its spectroscopic radial velocity (RV). Through a joint analysis of Betelgeuse’s brightness and RV variations and the phase offset between them compared to a sample of other stars with measured LSPs and RVs, Goldberg et al. (2024) argued that the binarity hypothesis is the best explanation for Betelgeuse’s LSP and RV signal, systematically ruling out all other proposed causes of the LSP in Betelgeuse. MacLeod et al. (2025) infer the existence of an using a long baseline of RV data as well as astrometric analysis, finding a periodic signal consistent with orbital motion on the LSP timescale in both data sets. In both of these works, the LSP is attributed to a low-mass () binary companion on a yr, AU, low-eccentricity orbit, at relatively low inclination relative to the observer. A companion at such a low mass and wide orbital separation would not directly affect Betelgeuse’s evolution through any substantial mass transfer, but it may interact indirectly via tides or by accreting from Betelgeuse’s stellar wind.
@@
The idea that Betelgeuse could presently be a binary system has been explored in the literature. For example, from speckle-imaging measurements Karovska et al. (1986) proposed that Betelgeuse has two high-eccentricity companions, including one at a separation of 0.06” (near that of ) but a period of 2.1 years.111This particular configuration was not confirmed in follow-up work with higher-resolution instruments (e.g. Wilson et al., 1992; Kervella et al., 2009; Montargès et al., 2016), nor recovered in the lightcurve or RV signal (Joyce et al., 2020; Goldberg et al., 2024; MacLeod et al., 2023, 2025). Moreover, a companion orbiting Betelgeuse is consistent with expectations for stars in Betelgeuse’s mass class (). Betelgeuse evolved from a late-O/early-B, main-sequence (MS) star. The binary fraction of late-O/early-B stars is 95% (e.g. Sana et al., 2012; de Mink et al., 2013). Though close binaries ( AU, which would have interacted by the RSG phase) show a uniform distribution of mass ratio with an excess at , at wider separations the mass ratio companions are skewed toward small mass ratios, favoring at 10AU (Rizzuto et al., 2013; Moe & Di Stefano, 2017; Offner et al., 2023). Other systems have been confirmed to have companions with low and intermediate separations, such as a large population of Cepheids in the range with binary companions on 1-20 yr periods, identified via both RV measurements (e.g. Evans et al., 2015) and X-ray emission and UV excess (Evans et al., 2022). It is therefore unsurprising that such a massive star has a companion at AU.
@@
Moving beyond dynamical arguments, observational follow-up informed by the companion’s orbital configuration will further constrain its nature and impact on the star and its environment. Importantly, Goldberg et al. (2024) show that the phase offset between the RV and light curves requires the companion to be behind Betelgeuse during Betelgeuse’s LSP brightness minima. The Great Dimming, which prompted many immediate follow-up observations, occurred at the brightness minimum of both the stellar pulsation and the LSP. However, at the time the companion was behind the disc of Betelgeuse, so observations then (e.g. Dupree et al., 2020, 2022; Kashyap et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; Montargès et al., 2021; Kravchenko et al., 2021; Alexeeva et al., 2021; Matthews & Dupree, 2022; Taniguchi et al., 2022) would not show any emission from the companion. Epochs corresponding to other orbital configurations, on the other hand, are ripe for further exploration.
@@
MacLeod et al. (2025) further argue that the companion could plausibly induce rotational velocities of km/s via tidal interactions, matching the measured 36 year period in the lightcurve and RV (Joyce et al., 2020; MacLeod et al., 2025) as well as observations of a dipolar “rotational” velocity field in spatially-resolved observations (redshifted in the SE and blueshifted in the NW, see, e.g. Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018). Fast rotation in Betelgeuse has previously been argued to indicate a recent merger, as rapid rotation of a RSG at the km/s level requires an external source of angular momentum input into the star’s outer envelope (e.g. Wheeler et al., 2017; Wheeler & Chatzopoulos, 2023; Chatzopoulos et al., 2020; Nance et al., 2018; Shiber et al., 2024).222See, however, discussion by Ma et al. (2024) on under-resolved large-scale convection as a plausible origin for the Kervella et al. (2018) velocity field, a counter-hypothesis which predicts a dipole orientation which changes stochastically and decoheres on a year-long convective overturn timescale. If the companion is instead responsible, then when combined with an orbital phase inferred from Betelgeuse’s RV, the direction of rotation (PA of 48 W. of S.; Kervella et al. 2018) would constrain the companion’s projected orbital plane and thus on-sky location relative to Betelgeuse.
@@
We obtained Director’s Discretionary Time (DDT) during Hubble Space Telescope (HST) Cycle 32 to observe Betelgeuse and its putative companion with the Space Telescope Imaging Spectrograph (STIS) instrument in the Far-UV (FUV) in a configuration never before observed. Though MacLeod et al. (2025) argue that the chromospheric emission from Betelgeuse would prevent detection of a companion, the mass range—which extends potentially up to permitted by the high-quality STELLA RVs (Goldberg et al., 2024) and consistent with the astrometric analysis of MacLeod et al. (2025)—includes a number of scenarios that could be detectable in both the far-UV and X-ray bands (see our companion publication in O’Grady et al. 2025).
@@ 2 Expectations for UV emission
Our observing strategy is informed by the various mass estimates and constraints for , and the most likely evolutionary possibilities for its companion. In this section we discuss the range of expectations for the companion’s spectral properties, and emphasize the importance of observing at ’s peak expected offset in redshift from Betelgeuse.
@@ 2.1 Possible Companion Scenarios
A mass range of 0.5–2 M⊙ enables a range of possible identities for the companion. One astrophysical object satisfying the mass condition is a low-mass, “standard” star. Under co-natal and coeval assumptions for the system (Dolan et al., 2016; Joyce et al., 2020; Neuhäuser et al., 2022), such a star must in fact be a young stellar object, or YSO, on the basis of Betelgeuse’s age of roughly 10 Myr (Tognelli et al., 2011; Baraffe et al., 2015; Haemmerlé et al., 2019; Amard et al., 2019). It is possible that the companion is not co-natal, but rather has been captured due to dynamical interactions, with the system then ejected from its birth cluster. However, Betelgeuse’s proper motion has not yet been traced back to a specific cluster (though see some discussion in Harper et al., 2008; Bouy et al., 2014; Bouy & Alves, 2015). Moreover, ejections generally happen within 1 Myr of cluster formation (Oh & Kroupa, 2016), but the age of Betelgeuse and all nearby clusters and OB associations are much older than 1 Myr (e.g. Blaauw, 1964; Briceño et al., 2007, 2019), so even a dynamical ejection scenario would be close to coeval.
Other possibilities include a more exotic, compact object, such as a neutron star (NS) or white dwarf (WD). In the case of a WD, it is not possible to construct a non-interacting, two-body, coeval evolutionary scenario that combines a 10-Myr-old red supergiant with even the highest-mass WD, which has a theoretical lower age limit of 40 Myr and a practical (observed) lower age limit of 70 Myr. While more complex three-body (or more) scenarios involving the deposition of a WD into Betelgeuse’s orbit from some other system are theoretically possible, the kinematic signatures of such an event are unclear. Evolutionary timescales therefore preclude the most likely versions of a WD scenario.
Additionally, in the case of a compact object, we would expect to see a bright X-ray signature of accretion: given the orbital separation, the companion is embedded within the dusty shell surrounding Betelgeuse (e.g., Haubois et al., 2019, 2023), and would interact with Betelgeuse’s wind (Aizu, 1973; Livio et al., 1986; Blondin et al., 1990; Mukai, 2017; Maldonado et al., 2025). Typical X-ray luminosities for accreting WDs are erg s-1 (Luna et al., 2013), and for accreting NSs are erg s-1 (Yungelson et al., 2019; De et al., 2022a, b). Complementary observations with Chandra were performed within a few weeks of the HST observations and revealed no such signature, as discussed in detail in O’Grady et al. (2025).